OpenStudy (anonymous):

What is the sum of a 10–term arithmetic sequence where the first term is 3 and the last term is 75?

OpenStudy (amistre64):

A = {3+...+7} +A = {7+...+3} -------------- 2A = {10+...+10} 2A = 10*10 = 100 A = 100/2 = 50

OpenStudy (amistre64):

opps, last term is 75 not 7 :)

OpenStudy (amistre64):

same idea tho (75+3)10/2 78*5